# Reinforcement Learning in Vision: A Survey

### v1

@@ 3.2.1 Image Generation
RL for image generation models departs from the language counterpart in both action space continuous diffusion steps or prompt refinements and reward design, which must capture perceptual quality, text–image alignment, and subject fidelity. A first family learns an explicit visual reward model: ImageReward (Xu et al., 2023) supplies human-preference scores that drive policy-gradient fine-tuning of diffusion backbones in DiffPPO (Xiao et al., 2024), Dpok (Fan et al., 2023a), and FocusDiff (Pan et al., 2025). A second line bypasses reward modelling by optimising pairwise or unary preferences: DDPO (Black et al., 2023), DiffusionDPO (Wallace et al., 2024), Diffusion-KTO (Li et al., 2024b), and DiffusionRPO (Gu et al., 2024) treat denoising trajectories as MDPs and apply R1/GRPO updates to maximise comparative human feedback. Beyond alignment, works such as PrefPaint (Liu et al., 2024a), Parrot (Lee et al., 2024), and RLRF (Rodriguez et al., 2025b) craft multi-objective or render-and-compare rewards to refine aesthetics, diversity, or vector graphics. RL has also been used to inject reasoning and prompt adaptation: ReasonGen-R1 (Zhang et al., 2025h), GoT-R1 (Duan et al., 2025) and RePrompt (Wu et al., 2025b) first generate textual plans or improved prompts, then reinforce the generator for coherent scene synthesis. Finally, personalisation methods—DPG-T2I (Wei et al., 2024), RPO (Miao et al., 2024a), and B2-DiffuRL (Hu et al., 2025a) optimize rewards that measure subject fidelity under scarce reference images. Collectively, these studies show that RL, armed with perceptual or preference-based rewards, can steer diffusion models toward higher realism, stronger prompt faithfulness, controllable layout, and user-specific appearance capabilities difficult to achieve with likelihood training alone.
@@ 3.2.2 Video Generation
Applying RL to video generation introduces challenges absent in images: rewards must capture temporal coherence, motion naturalness, and text–video alignment across hundreds of frames. Early work such as InstructVideo (Yuan et al., 2024) repurposed image scorers and applied PPO to refine short clips, whereas VideoRM (Wu et al., 2024) and VideoReward (Liu et al., 2025a) learn dedicated preference models that grade entire sequences for smoothness, consistency and prompt faithfulness. Building on GRPO/R1, DanceGRPO (Xue et al., 2025) shows that group-normalized returns stabilize long-horizon optimization and boost aesthetic and alignment scores across diffusion and flow backbones.
@@
A complementary line adopts direct preference optimization. DreamDPO (Zhou et al., 2025c) extends DPO to NeRF and mesh diffusion models by sampling paired 3D outputs and maximizing the margin dictated by human comparisons, achieving superior prompt fidelity without an explicit value network. Finally, Nabla-R2D3 (Liu et al., 2025c) tackles interactive scene editing: the agent sequentially adds, deletes, or transforms objects in a 3D scene; reward is computed via real-time rasterized views and task-specific validators (e.g., occupancy, spatial relations). Group-relative policy optimization (R2D3) stabilizes training despite sparse feedback. Together, these studies demonstrate that RL, equipped with geometry-aware or render-based rewards, provides an effective mechanism for controlling structural integrity, text alignment, and interactive editability capabilities that standard likelihood or score-distillation objectives struggle to capture in the 3D domain.
@@ 3.3.1 Unified RL
UniRL (Mao et al., 2025) exemplifies this paradigm: a visual autoregressive backbone is first instruction-tuned, then jointly fine-tuned on VQA, captioning and image generation with a blended reward measuring textual correctness, CLIP-based alignment, and aesthetic quality. CoRL (Jiang et al., 2025b) pushes the idea further by alternating “co-understanding” and “co-generation” batches within the same GRPO step. To address inefficiency in dense token spaces, SelfTok (Wang et al., 2025a) discretises multi-modal actions into a self-evolving token set and demonstrates that a single RL head can govern retrieval, grounding, and synthesis with minimal extra parameters. Finally, HermesFlow (Yang et al., 2025b) couples an autoregressive text module with a rectified flow image decoder under one cross-task reward, illustrating that diffusion-style and language-style policies can be harmonized through unified reinforcement updates. Together, these works suggest that sharing a common RL objective across tasks not only reduces training cost but also encourages emergent cross modal generalization unavailable to isolated, task specific fine-tuning.
@@ 3.4.1 GUI Automation
Modern GUI RL research frames screen understanding and action prediction as a vision–language decision process, then employs rule-based or preference rewards to close the perception–action loop. On desktop and web interfaces, GUI-R1 (Luo et al., 2025) introduces an R1-style rule set that maps click success, text entry, and scroll validity to dense rewards. UI-R1 (Lu et al., 2025) adds GRPO with a novel action-specific KL term to stabilize long-horizon plans, while SE-GUI (Yuan et al., 2025b) applies self-evolutionary filtering to distil high-fidelity trajectories. Focusing on trajectory reuse, UIShift (Gao et al., 2025b) formulates an inverse dynamics objective that lets MLLM learn actions from unlabeled GUI pairs and then refines them via RL. Complementary preference-based frameworks include LPO (Tang et al., 2025) that rewards spatial proximity for precise clicks. ProgRM (Zhang et al., 2025a) injects program-level logical checks, and RUIG (Zhang et al., 2023) leverages instruction grounding with reinforcement signals. Tool-specific baselines such as Ui-tars (Qin et al., 2025) offer larger action vocabularies yet still rely on rule-driven RL for robust execution.
Mobile scenarios introduce latency and on-device constraints. AgentCPM-GUI (Zhang et al., 2025i) compresses the action space and conducts GRPO fine-tuning. MobileGUI-RL (Shi et al., 2025) advances this line via online RL with task-level rewards to improve exploration under limited memory, and Mobile-R1 (Gu et al., 2025) extends interactive multi-turn RL to correct error cascades during long tasks. At inference, GTA1 (Yang et al., 2025d) samples multiple action candidates and employs a judge model to pick the best, effectively trading compute for higher success rates. Additional, light-weight models such as Appvlm (Papoudakis et al., 2025) demonstrate that modest-sized MLLM, after GRPO fine-tuning, can control smartphone apps with competitive precision. To adaptively reduce the thinking length, the TON (Wang et al., 2025b) proposes a thought-dropout solution during supervised fine-tuning stage, then GRPO skill adaptivley skip unnecessary reasoning process for efficiently thinking.
@@ 3.4.3 Visual Manipulation
Visual manipulation tasks (e.g.,  object relocation, tool use, and multi-step rearrangement) require fine-grained perception and long-horizon planning. Recent works (Chen et al., 2025g; Shu et al., 2025) integrate reinforcement learning with vision–language models to enhance generalization, interactivity, and policy consistency. TGRPO (Chen et al., 2025g) introduces a task-grounded reward formulation and group normalized updates to stabilize training for open-ended object manipulation. RFTF (Shu et al., 2025) applies rule-based rewards to support interactive table top tasks and emphasizes training with minimal human supervision. Meanwhile, RLVLA (Liu et al., 2025b) and VLA-RL (Lu et al., 2025b) explore curriculum-based or progressive reinforcement learning for VLM-based robot agents, achieving high success rates across diverse rearrangement environments.
Building on this, ConRFT (Chen et al., 2025e) and iRe-VLA (Guo et al., 2025b) introduce consistency-aware and instruction-refinement strategies respectively, using RL to align visual predictions with physical interaction outcomes. RIPT-VLA focuses on interactive prompting during manipulation, bridging LLM planning and low-level control through reinforced feedback (Tan et al., 2025b). Finally, ReinBot (Zhang et al., 2025b) leverages multimodal rollouts and preference-based updates to improve real-world manipulation robustness (Zhang et al., 2025b). Collectively, these studies highlight the role of vision-language reasoning, structured reward design, and RL-based refinement in advancing embodied manipulation under complex, language-conditioned settings.
@@
Verifiable Rewards: Deepseekmath (Shao et al., 2024b), IoU (Rezatofighi et al., 2019), CodeRL (Le et al., 2022).
@@
Reward from Human Preferencee.g., ImageReward (Xu et al., 2023), HPS (Wu et al., 2023b), HPS V2 (Wu et al., 2023b), Pick-a-pic (Kirstain et al., 2023), VideoReward (Liu et al., 2025a), RichHF-18K (Liang et al., 2024).
@@
Reward from Model Preferencee.g., VideoPrefer (Wu et al., 2024), PARM (Zhang et al., 2025e).
Denoising Trajectory Diagnostics: e.g.,  VARD (Dai et al., 2025), Inversion-DPO (Li et al., 2025c).
@@
Unified Models Generation–oriented Task Metricse.g., GenEval (Ghosh et al., 2023), DPG-Bench (Hu et al., 2024b), ParaPrompts (Wu et al., 2025c).
@@
KL-divergencee.g., HermesFlow (Yang et al., 2025b)
@@
GUI Action Accuracy (offline)e.g., ScreenSpot (Cheng et al., 2024), ScreenSpot-Pro (Li et al., 2025a), Ui-vision (Nayak et al., 2025).
@@
Model Preference Critic: e.g., ProgRM (Zhang et al., 2025a).
@@ 4.3 Evaluation of Visual Generation Models
Reward Metric. During RL fine-tuning, each generated image or video receives a sample-level reward $\mathcal{M}_{\mathrm{samp}}$. Human-preference rewards, e.g., ImageReward (Xu et al., 2023) and HPS (Wu et al., 2023b) supply dense signals that correlate well with aesthetic appeal. When a deterministic checker exists, authors turn to verifiable rewards: MotionPrompt (Nam et al., 2025) and DSPO (Cai et al., 2025b) use optical-flow, object masks that can be evaluated without humans. A third route relies on model preference rewards, where a stronger frozen critic (e.g., VideoPrefer (Wu et al., 2024) or PARM (Zhang et al., 2025e)) scores samples, enabling scalable DPO/PPO training.
Model-State Metric. Two light diagnostics track training stability. (i) Denoising trajectory statistics: VARD (Dai et al., 2025) and Inversion DPO (Li et al., 2025c) record per-step noise predictions or DDIM traces; pathological spikes reveal early collapse. (ii) KL divergence between the current diffusion policy and its frozen base ($\pi_{\theta_{t}}\!\parallel\!\pi_{\theta_{0}}$), popularized by DDPO (Black et al., 2023) and reused in VARD.
@@ 4.4 Evaluation of Unified Models
Task Metric. Two benchmark families are widely adopted. Generation-oriented suites such as GenEval (Ghosh et al., 2023), DPG-Bench (Hu et al., 2024b), and ParaPrompts (Wu et al., 2025c) focus on prompt-faithful generation, testing multi-object composition, style control, and long-caption adherence. Conversely, understanding-oriented benchmarks (MME (Fu et al., 2023), POPE (Li et al., 2023)) measure grounding, reasoning and hallucination detection from the same backbone.
@@
Model-State Metric. Unified models additionally track fine-grained diagnostics during RL. UniRL (Mao et al., 2025) proposes a generation–understanding imbalance score the absolute gap between batch-level rewards on the two task families to prevent one modality from dominating the update. HermesFlow (Yang et al., 2025b) monitors the KL divergence between the current shared policy $\pi_{\theta_{t}}$ and its supervised baseline $\pi_{\theta_{0}}$ on both generation and understanding prompts, serving as an early-warning signal for policy collapse. These state-level curves $\mathcal{M}_{\mathrm{state}}^{t}$ allow practitioners to stop or re-weight training before external task scores degrade.
@@ 4.5 Evaluation of Vision Language Action Models
Time-R1 (Wang et al., 2025g) Mar 2025 Tr&Te 2.5 K TimeRFT grounding spans (IoU reward)
@@
IntentBench (Yang et al., 2025c) Jun 2025 Te Omnimodal evaluation demands unified audio-visual comprehension
@@
VideoReward (Liu et al., 2025a) Jan 2025 Te Human-ranked video pairs across quality, motion, and text alignment (used in (Liu et al., 2025a))
@@
CAGUI (Zhang et al., 2025i) Jun 2025 Te 55 K trajectories from 30 Chinese Android apps, 8 domains
@@
GUICours (Chen et al., 2024) Jun 2024 Tr&Te GUIEnv/Act/Chat datasets (10 M OCR + 67 k navigation)
@@ 4.6 Benchmarks
Benchmarks for visual generation tasks predominantly supply human preference data that serve as reward models for policy optimization. Datasets like ImageReward (Xu et al., 2023) and HPS v1 & v2 (Wu et al., 2023b; b) collect human-ranked pairs of text-to-image outputs, allowing one to train a scalar reward function that scores generations. Such reward models have been used to refine text-to-image diffusion models via RLHF, aligning outputs with human aesthetic preferences. Similarly, Pick-a-Pic (Kirstain et al., 2023) and VideoReward  (Liu et al., 2025a)extend this to broader user preferences (motion smoothness, text alignment). Some benchmarks also facilitate robust evaluation of generative RL agents on generalization. T2I-CompBench (Huang et al., 2023) is a text-to-image compositionality test set that requires correctly binding novel combinations of attributes and object relations – a measure of compositional generalization often used to assess RL-trained generators. Likewise, domain-specific benchmarks define verifiable success criteria as rewards: StarVector (Rodriguez et al., 2025a) provides SVG code-generation tasks with a strict shape-matching reward, and AnimeReward (Zhu et al., 2025a) targets consistency in animated video generations with multi-dimensional human preference scores (image-video coherence, character consistency, etc.).
@@
For vision–language action agents, numerous benchmarks provide expert trajectories and simulated environments with clear reward signals for policy training and robust evaluation. Many are centered on GUI and web interaction tasks, where success can be unambiguously measured. For instance, GUI-R1-3K (Luo et al., 2025) compiles $3,000+$ GUI manipulation trajectories across Windows, Linux, macOS, Android, and web platforms. It introduces an “R1-style” dense reward scheme mapping each correct action (e.g., clicking the right button, entering correct text, a valid scroll) to positive feedback, providing step-wise reinforcement that guides an agent through multi-step UI tasks. Building on this, SE-GUI (Yuan et al., 2025b) curates 3k high-quality GUI examples with grounded instructions and bounding-box annotations, which are used to train agents with a self-imitation RL strategy. Evaluation-focused suites like UI-R1 (Lu et al., 2025) define a fixed set of unseen tasks (e.g.,   136 mobile GUI tasks spanning click, scroll, swipe, text-input actions) to test generalization of learned policies. Meanwhile, web interaction benchmarks such as Mind2Web (Deng et al., 2023) offer  2,000 tasks on real websites with a binary success/failure reward for completing each task. Some datasets emphasize exact match and reproducibility: AITZ (Zhang et al., 2024b) (Android Interaction w/ CoAT reasoning) logs  18k screen-action pairs with corresponding tool-assisted rationales, and uses an exact action match reward to ensure precise adherence to instructions. On the other hand, broader benchmarks like OmniAct (Kapoor et al., 2024) and GUICoURS (Chen et al., 2024) target generalist agent capabilities across diverse domains. OmniAct integrates nearly 10k scripted desktop and web tasks into a single environment, while GUICoURS combines multimodal resources (10M OCR observations, 67k navigation demonstrations) spanning GUI, web, and chat interfaces. The reward structures in these benchmarks are carefully crafted, from rule-based metrics to preference scores, to guide policy learning and reflect task goals. They enable visual RL agents to learn from meaningful feedback and be evaluated not just on task success, but also on alignment with human reasoning and performance on complex, long-horizon decisions.

### v2

@@ 3.2.1 Image Generation
RL for image generation models departs from the language counterpart in both action space continuous diffusion steps or prompt refinements and reward design, which must capture perceptual quality, text–image alignment, and subject fidelity. A first family learns an explicit visual reward model: ImageReward (Xu et al., 2023) supplies human-preference scores that drive policy-gradient fine-tuning of diffusion backbones in DiffPPO (Xiao et al., 2024), Dpok (Fan et al., 2023a), and FocusDiff (Pan et al., 2025). A second line bypasses reward modelling by optimising pairwise or unary preferences: DDPO (Black et al., 2023), DiffusionDPO (Wallace et al., 2024), Diffusion-KTO (Li et al., 2024b), and DiffusionRPO (Gu et al., 2024) treat denoising trajectories as MDPs and apply R1/GRPO updates to maximise comparative human feedback. Beyond alignment, works such as PrefPaint (Liu et al., 2024a), Parrot (Lee et al., 2024), and RLRF (Rodriguez et al., 2025b) craft multi-objective or render-and-compare rewards to refine aesthetics, diversity, or vector graphics. RL has also been used to inject reasoning and prompt adaptation: ReasonGen-R1 (Zhang et al., 2025i), GoT-R1 (Duan et al., 2025) and RePrompt (Wu et al., 2025b) first generate textual plans or improved prompts, then reinforce the generator for coherent scene synthesis. Finally, personalisation methods—DPG-T2I (Wei et al., 2024), RPO (Miao et al., 2024a), and B2-DiffuRL (Hu et al., 2025a) optimize rewards that measure subject fidelity under scarce reference images. Collectively, these studies show that RL, armed with perceptual or preference-based rewards, can steer diffusion models toward higher realism, stronger prompt faithfulness, controllable layout, and user-specific appearance capabilities difficult to achieve with likelihood training alone.
@@ 3.2.2 Video Generation
Applying RL to video generation introduces challenges absent in images: rewards must capture temporal coherence, motion naturalness, and text–video alignment across hundreds of frames. Early work such as InstructVideo (Yuan et al., 2024) repurposed image scorers and applied PPO to refine short clips, whereas VideoRM (Wu et al., 2024) and VideoReward (Liu et al., 2025b) learn dedicated preference models that grade entire sequences for smoothness, consistency and prompt faithfulness. Building on GRPO/R1, DanceGRPO (Xue et al., 2025) shows that group-normalized returns stabilize long-horizon optimization and boost aesthetic and alignment scores across diffusion and flow backbones.
@@
A complementary line adopts direct preference optimization. DreamDPO (Zhou et al., 2025c) extends DPO to NeRF and mesh diffusion models by sampling paired 3D outputs and maximizing the margin dictated by human comparisons, achieving superior prompt fidelity without an explicit value network. Finally, Nabla-R2D3 (Liu et al., 2025d) tackles interactive scene editing: the agent sequentially adds, deletes, or transforms objects in a 3D scene; reward is computed via real-time rasterized views and task-specific validators (e.g., occupancy, spatial relations). Group-relative policy optimization (R2D3) stabilizes training despite sparse feedback. Together, these studies demonstrate that RL, equipped with geometry-aware or render-based rewards, provides an effective mechanism for controlling structural integrity, text alignment, and interactive editability capabilities that standard likelihood or score-distillation objectives struggle to capture in the 3D domain.
@@ 3.3.1 Unified RL
UniRL (Mao et al., 2025) exemplifies this paradigm: a visual autoregressive backbone is first instruction-tuned, then jointly fine-tuned on VQA, captioning and image generation with a blended reward measuring textual correctness, CLIP-based alignment, and aesthetic quality. CoRL (Jiang et al., 2025b) pushes the idea further by alternating “co-understanding” and “co-generation” batches within the same GRPO step. To address inefficiency in dense token spaces, SelfTok (Wang et al., 2025a) discretises multi-modal actions into a self-evolving token set and demonstrates that a single RL head can govern retrieval, grounding, and synthesis with minimal extra parameters. Finally, HermesFlow (Yang et al., 2025c) couples an autoregressive text module with a rectified flow image decoder under one cross-task reward, illustrating that diffusion-style and language-style policies can be harmonized through unified reinforcement updates. Together, these works suggest that sharing a common RL objective across tasks not only reduces training cost but also encourages emergent cross modal generalization unavailable to isolated, task specific fine-tuning.
@@ 3.4.1 GUI Automation
Modern GUI RL research frames screen understanding and action prediction as a vision–language decision process, then employs rule-based or preference rewards to close the perception–action loop. On desktop and web interfaces, GUI-R1 (Luo et al., 2025) introduces an R1-style rule set that maps click success, text entry, and scroll validity to dense rewards. UI-R1 (Lu et al., 2025) adds GRPO with a novel action-specific KL term to stabilize long-horizon plans, while SE-GUI (Yuan et al., 2025b) applies self-evolutionary filtering to distil high-fidelity trajectories. Focusing on trajectory reuse, UIShift (Gao et al., 2025b) formulates an inverse dynamics objective that lets MLLM learn actions from unlabeled GUI pairs and then refines them via RL. Complementary preference-based frameworks include LPO (Tang et al., 2025) that rewards spatial proximity for precise clicks. ProgRM (Zhang et al., 2025b) injects program-level logical checks, and RUIG (Zhang et al., 2023) leverages instruction grounding with reinforcement signals. Tool-specific baselines such as Ui-tars (Qin et al., 2025) offer larger action vocabularies yet still rely on rule-driven RL for robust execution.
Mobile scenarios introduce latency and on-device constraints. AgentCPM-GUI (Zhang et al., 2025j) compresses the action space and conducts GRPO fine-tuning. MobileGUI-RL (Shi et al., 2025) advances this line via online RL with task-level rewards to improve exploration under limited memory, and Mobile-R1 (Gu et al., 2025) extends interactive multi-turn RL to correct error cascades during long tasks. At inference, GTA1 (Yang et al., 2025e) samples multiple action candidates and employs a judge model to pick the best, effectively trading compute for higher success rates. Additional, light-weight models such as Appvlm (Papoudakis et al., 2025) demonstrate that modest-sized MLLM, after GRPO fine-tuning, can control smartphone apps with competitive precision. To adaptively reduce the thinking length, the TON (Wang et al., 2025b) proposes a thought-dropout solution during supervised fine-tuning stage, then GRPO skill adaptivley skip unnecessary reasoning process for efficiently thinking.
@@ 3.4.3 Visual Manipulation
Visual manipulation tasks (e.g.,  object relocation, tool use, and multi-step rearrangement) require fine-grained perception and long-horizon planning. Recent works (Chen et al., 2025g; Shu et al., 2025) integrate reinforcement learning with vision–language models to enhance generalization, interactivity, and policy consistency. TGRPO (Chen et al., 2025g) introduces a task-grounded reward formulation and group normalized updates to stabilize training for open-ended object manipulation. RFTF (Shu et al., 2025) applies rule-based rewards to support interactive table top tasks and emphasizes training with minimal human supervision. Meanwhile, RLVLA (Liu et al., 2025c) and VLA-RL (Lu et al., 2025b) explore curriculum-based or progressive reinforcement learning for VLM-based robot agents, achieving high success rates across diverse rearrangement environments.
Building on this, ConRFT (Chen et al., 2025e) and iRe-VLA (Guo et al., 2025b) introduce consistency-aware and instruction-refinement strategies respectively, using RL to align visual predictions with physical interaction outcomes. RIPT-VLA focuses on interactive prompting during manipulation, bridging LLM planning and low-level control through reinforced feedback (Tan et al., 2025b). Finally, ReinBot (Zhang et al., 2025c) leverages multimodal rollouts and preference-based updates to improve real-world manipulation robustness (Zhang et al., 2025c). Collectively, these studies highlight the role of vision-language reasoning, structured reward design, and RL-based refinement in advancing embodied manipulation under complex, language-conditioned settings.
@@
Verifiable Rewards: e.g.,  Deepseekmath (Shao et al., 2024b), IoU (Rezatofighi et al., 2019), CodeRL (Le et al., 2022).
@@
Reward from Human Preferencee.g., ImageReward (Xu et al., 2023), HPS (Wu et al., 2023b), HPS V2 (Wu et al., 2023b), Pick-a-pic (Kirstain et al., 2023), VideoReward (Liu et al., 2025b), RichHF-18K (Liang et al., 2024).
@@
Reward from Model Preferencee.g., VideoPrefer (Wu et al., 2024), PARM (Zhang et al., 2025f).
Denoising Trajectory Diagnostics: e.g.,  VARD (Dai et al., 2025), Inversion-DPO (Li et al., 2025f).
@@
Unified Models Generation–oriented Task Metricse.g., GenEval (Ghosh et al., 2023), DPG-Bench (Hu et al., 2024b), ParaPrompts (Wu et al., 2025d).
@@
KL-divergencee.g., HermesFlow (Yang et al., 2025c).
@@
GUI Action Accuracy (offline)e.g., ScreenSpot (Cheng et al., 2024), ScreenSpot-Pro (Li et al., 2025c), Ui-vision (Nayak et al., 2025).
@@
Model Preference Critic: e.g., ProgRM (Zhang et al., 2025b).
@@ 4.3 Evaluation of Visual Generation Models
Reward Metric. During RL fine-tuning, each generated image or video receives a sample-level reward $\mathcal{M}_{\mathrm{samp}}$. Human-preference rewards, e.g., ImageReward (Xu et al., 2023) and HPS (Wu et al., 2023b) supply dense signals that correlate well with aesthetic appeal. When a deterministic checker exists, authors turn to verifiable rewards: MotionPrompt (Nam et al., 2025) and DSPO (Cai et al., 2025b) use optical-flow, object masks that can be evaluated without humans. A third route relies on model preference rewards, where a stronger frozen critic (e.g., VideoPrefer (Wu et al., 2024) or PARM (Zhang et al., 2025f)) scores samples, enabling scalable DPO/PPO training.
Model-State Metric. Two light diagnostics track training stability. (i) Denoising trajectory statistics: VARD (Dai et al., 2025) and Inversion DPO (Li et al., 2025f) record per-step noise predictions or DDIM traces; pathological spikes reveal early collapse. (ii) KL divergence between the current diffusion policy and its frozen base ($\pi_{\theta_{t}}\!\parallel\!\pi_{\theta_{0}}$), popularized by DDPO (Black et al., 2023) and reused in VARD.
@@ 4.4 Evaluation of Unified Models
Task Metric. Two benchmark families are widely adopted. Generation-oriented suites such as GenEval (Ghosh et al., 2023), DPG-Bench (Hu et al., 2024b), and ParaPrompts (Wu et al., 2025d) focus on prompt-faithful generation, testing multi-object composition, style control, and long-caption adherence. Conversely, understanding-oriented benchmarks (MME (Fu et al., 2023), POPE (Li et al., 2023)) measure grounding, reasoning and hallucination detection from the same backbone.
@@
Model-State Metric. Unified models additionally track fine-grained diagnostics during RL. UniRL (Mao et al., 2025) proposes a generation–understanding imbalance score the absolute gap between batch-level rewards on the two task families to prevent one modality from dominating the update. HermesFlow (Yang et al., 2025c) monitors the KL divergence between the current shared policy $\pi_{\theta_{t}}$ and its supervised baseline $\pi_{\theta_{0}}$ on both generation and understanding prompts, serving as an early-warning signal for policy collapse. These state-level curves $\mathcal{M}_{\mathrm{state}}^{t}$ allow practitioners to stop or re-weight training before external task scores degrade.
@@ 4.5 Evaluation of Vision Language Action Models
Time-R1 (Wang et al., 2025h) Mar 2025 Tr&Te 2.5 K TimeRFT grounding spans (IoU reward)
@@
IntentBench (Yang et al., 2025d) Jun 2025 Te Omnimodal evaluation demands unified audio-visual comprehension
@@
VideoReward (Liu et al., 2025b) Jan 2025 Te Human-ranked video pairs across quality, motion, and text alignment (used in (Liu et al., 2025b))
@@
CAGUI (Zhang et al., 2025j) Jun 2025 Te 55 K trajectories from 30 Chinese Android apps, 8 domains
@@
GUICours (Chen et al., 2024b) Jun 2024 Tr&Te GUIEnv/Act/Chat datasets (10 M OCR + 67 k navigation)
@@ 4.6 Benchmarks
Benchmarks for visual generation tasks predominantly supply human preference data that serve as reward models for policy optimization. Datasets like ImageReward (Xu et al., 2023) and HPS v1 & v2 (Wu et al., 2023b; b) collect human-ranked pairs of text-to-image outputs, allowing one to train a scalar reward function that scores generations. Such reward models have been used to refine text-to-image diffusion models via RLHF, aligning outputs with human aesthetic preferences. Similarly, Pick-a-Pic (Kirstain et al., 2023) and VideoReward  (Liu et al., 2025b)extend this to broader user preferences (motion smoothness, text alignment). Some benchmarks also facilitate robust evaluation of generative RL agents on generalization. T2I-CompBench (Huang et al., 2023) is a text-to-image compositionality test set that requires correctly binding novel combinations of attributes and object relations – a measure of compositional generalization often used to assess RL-trained generators. Likewise, domain-specific benchmarks define verifiable success criteria as rewards: StarVector (Rodriguez et al., 2025a) provides SVG code-generation tasks with a strict shape-matching reward, and AnimeReward (Zhu et al., 2025a) targets consistency in animated video generations with multi-dimensional human preference scores (image-video coherence, character consistency, etc.).
@@
For vision–language action agents, numerous benchmarks provide expert trajectories and simulated environments with clear reward signals for policy training and robust evaluation. Many are centered on GUI and web interaction tasks, where success can be unambiguously measured. For instance, GUI-R1-3K (Luo et al., 2025) compiles $3,000+$ GUI manipulation trajectories across Windows, Linux, macOS, Android, and web platforms. It introduces an “R1-style” dense reward scheme mapping each correct action (e.g., clicking the right button, entering correct text, a valid scroll) to positive feedback, providing step-wise reinforcement that guides an agent through multi-step UI tasks. Building on this, SE-GUI (Yuan et al., 2025b) curates 3k high-quality GUI examples with grounded instructions and bounding-box annotations, which are used to train agents with a self-imitation RL strategy. Evaluation-focused suites like UI-R1 (Lu et al., 2025) define a fixed set of unseen tasks (e.g.,   136 mobile GUI tasks spanning click, scroll, swipe, text-input actions) to test generalization of learned policies. Meanwhile, web interaction benchmarks such as Mind2Web (Deng et al., 2023) offer  2,000 tasks on real websites with a binary success/failure reward for completing each task. Some datasets emphasize exact match and reproducibility: AITZ (Zhang et al., 2024b) (Android Interaction w/ CoAT reasoning) logs  18k screen-action pairs with corresponding tool-assisted rationales, and uses an exact action match reward to ensure precise adherence to instructions. On the other hand, broader benchmarks like OmniAct (Kapoor et al., 2024) and GUICoURS (Chen et al., 2024b) target generalist agent capabilities across diverse domains. OmniAct integrates nearly 10k scripted desktop and web tasks into a single environment, while GUICoURS combines multimodal resources (10M OCR observations, 67k navigation demonstrations) spanning GUI, web, and chat interfaces. The reward structures in these benchmarks are carefully crafted, from rule-based metrics to preference scores, to guide policy learning and reflect task goals. They enable visual RL agents to learn from meaningful feedback and be evaluated not just on task success, but also on alignment with human reasoning and performance on complex, long-horizon decisions.